# From Acceleration to Saturation: Scaling Behavior of Bootstrapped Language Model Pretraining

### v1

@@ B.1 Megatron-LM Configuration
Infrastructure. Our experiments are performed on multiple nodes, each consisting of 8 NVIDIA H100 80 GB GPUs, interconnected via InfiniBand HDR. The software we use for training is the Megatron-LM library [57].
We use and modify the Megatron-LM (core v0.8.0) library for our experiments [^4] https://github.com/NVIDIA/Megatron-LM. Models are trained with data type bfloat16. Other optimization libraries used include FlashAttention [16] and TransformerEngine [^5] https://github.com/NVIDIA/TransformerEngine. See the example scripts provided in supplementary material.
@@ B.2 Model Configuration
Let us elaborate more on our model configuration. The intermediate hidden dimension size, $d_{\rm MLP}$, is set to be four times the hidden dimension size, i.e., $4d_{\rm model}$. Bias is not used in the linear layers. We do not consider efficiency-motivated implementations like grouped query attention as well. Attention head number is chosen to increase with model size following standard practices. Other designs of the architecture follow Llama2’s closely [63].
We vary model sizes keeping the ratio $n_{\rm layer}/d_{\rm model}$ to lie in the range 32 to 64, as in [37]. Smaller models for used for ablation studies. See Table 3 for the exact numbers for the model configuration.
@@ B.3 Training Configuration
As discussed in the main text, we adopt the WSD learning rate schedule for all experiments. The number of warmup steps is set approximately equal to the model size, as suggested by [49]. In the final phase of training, the learning rate decays linearly to 10% of its peak value, with the decay phase spanning roughly 10% of the total training steps, following the setup in [27]. To emulate varying token budgets, we save intermediate checkpoints at logarithmically spaced intervals.
The general training configuration is summarized in Table 4, while model-specific hyperparameters, such as warmup iterations, initialization standard deviation ($\sqrt{2/5d_{\rm model}}$ [39]), maximum training steps, batch size, and tuned learning rate, are provided in Table 5. Batch size is scaled with model size according to standard practice, without tuning for optimality. The number of tokens is increased with model size as well, such that the parameter-to-token ratio remains roughly constant across different model sizes. 6 We note that no specialized techniques for mitigating training instabilities, such as Z-loss or QK normalization, are employed, as such instabilities do not arise in our experiments.
@@
Table 5: Model-dependent training configuration. "init. size" refers to the standard deviation of the normal distribution used for initializing the weights. "Max iter." refers to the maximum iteration run on the model.
Model warmup iter. init. size Max iter. batch size LR
@@
B.4 GPU Hours and Costs
@@
The estimate is as follows. We calculate the FLOPs for training the largest first and second-stage models with maximum iterations using the $6ND$ approximation, ignoring the additional FLOPs required to continued pretrain models with shorter iterations (as we can reuse the intermediate checkpoints). We further assume that the per-second TFLOPs of the GPU is 400. We obtain around $10^{9}$ TFLOPs. Taking into account of additional experimentation and ablation runs, we estimate that around 3000 GPU hours were used for the entire project. Assuming a cost of 2 USD per GPU hour, the total cost is around 6000 USD.
@@ C.1 Ablation of Learning Rate Schedules
Here, we compare the performances of using WSD and the commonly used learning rate (LR) cosine schedules (decaying to 10% of the peak LR value) [63]. The model size we use is 0.1B, with the following training configuration: batch size 512, 4,000 training iterations, and 200 warmup iterations. We can see from Figure 5 that both schedules yield similar (with WSD achieving slightly better final loss) performances. This justifies our choice of using the WSD schedule throughout the paper.
@@ C.3 Evaluation with Standard Benchmarks
We compare the performance of our trained 1B base model against existing models with similar sizes, Pythia [9] and TinyLlama [73], based on standard natural language processing benchmarks, ARC [14], lambada [46], logiqa [42], piqa [10], sciq [68], and winogrande [52].
@@
Table 6: Benchmarks’ performance comparison across models. Reported scores are accuracies (normalized by byte length whenever applicable). The first two columns are scores of existing models. The last column is evaluation results of the largest base model trained with the most number of tokens in this work. Our models are evaluated with the LM Evaluation Harness v0.4.0 library [22].
@@ C.4 Continual Pretraining Details
Using a lower LR for CPT is also a common practice [34]. As an ablation, we have experimented with using a constant LR schedule setting the value to be the final LR for the first-stage base model training, on a 0.1B model. We find that it yields worse performance (see Figure 7). Hence, we use the same LR (and LR schedule) as the base model for CPT.
@@ C.5 Model Growth Details
Stacking details. We first note that stacking does not preserve function but is empirically found to work well in practice [17]. We use the recommended stacking procedure in [17], which is, letting $M$ be the non-embedding part of the base model, the stacked model with growth factor $k$ is given by:
@@
Figure 7: Bootstrapped pretraining with lower LR. We show that using the same LR as the base model achieves better final performance than using a lower LR for a 0.1B model continual pretrained on code data (left) and stacking (right).
@@ D.1 Fitted Exponents and Other Results
We show the fitted exponents of the multiplicative scaling laws studied in Table 1 and 2 in Table 7.
@@
Figure 8: Left: $D_{1}$ has power-law scaling. We show scaling behavior of fist-stage training tokens ($D_{1}$) for different values of second-stage tokens ($D_{2}$), indicating that $D_{1}$ also has power-law scaling. From left to right: Continual pretraining on mathematics data, and model growth by expansion.
@@
D.2 Joint Scaling Law
@@
Figure 9: Fits of the joint upcycling scaling law of stacking. 10% of the collected data points with lowest losses are used for validation (orange points).
We provide the fitted coefficients of the joint scaling law in Table 8. In addition to bootstrapped pretraining, we also show the fitted coefficients for base models trained from scratch on the same dataset as the second stage (code data for CPT, and the same data as the base model for model growth). For model growth, we consider a growth factor of 2, and fit the parameters with model size $N$ before stacking for comparison conveniences. We further note that the fitted coefficients are produced by fitting the joint scaling law to all data points collected, including those used for validation in Figure 9.
@@
D.3 Why Fitting Scaling Laws Separately?
We justify our decision to fit two separate scaling laws—one for models trained from scratch and another for bootstrapped pretraining—instead of employing a unified formulation that spans all stages.
First, the scaling behavior of models trained from scratch is expected to differ from that of models undergoing bootstrapping. Specifically, at the limit $D_{1}\to 0$, grown models are initialized from a pretrained base model, whereas scratch-trained models start from random weights. This difference in initialization leads to distinct learning dynamics and, consequently, different scaling law parameters.
Similarly, the scaling behavior of the base model differs from that of the second-stage training in the limit $D_{2}\to 0$. As shown in Figure 7, second-stage training often begins with a rewarming phase, during which the loss initially increases before decreasing. This early instability deviates from the expected scaling of dense models, although the overall trend remains correlated—supporting the validity of Condition 1, due to function preservation and empirical observations that rewarming does not entirely disrupt loss behavior (as quoted from [15]: “In practice, few empirical phenomena obey power laws for all values of xx”).
@@ D.3 Why Fitting Scaling Laws Separately?
Appendix E More Practical Implications
To further examine the trade-off between training from scratch and bootstrapped pretraining, we fit Chinchilla-style scaling laws to models trained from scratch on the same datasets used in our second-stage experiments, namely, code or math data for CPT, and the original dataset for model growth. In Figure 10 and 11, we plot the validation loss as a function of training tokens for both approaches, across a range of sunk costs, using a fixed model size of 7B. For CPT on code data, we find that models trained from scratch begin to outperform bootstrapped models (pretrained on 10T tokens) once more than 300B tokens of domain-specific data are used.
Figure 10: Tokens required for from-scratch training to catch up with bootstrapped pretraining. We compare loss-versus-token plots of from-scratch and bootstrapped pretraining at various base model training budgets (sunk costs). Stacking is only considered efficient when $D_{\rm scratch}<$ sunk cost. We observe that the efficiency of stacking diminishes with sunk cost.
Figure 11: Tokens required for from-scratch training to catch up with bootstrapped pretraining. Conclusion similar to those made in Figure 10 can be made for CPT on code data, where the scaling efficiency decreases with first-stage training tokens.
E.1 Compute Optimality of Model Growth

### v2

@@ B.1 Megatron-LM Configuration
Infrastructure. Our experiments are performed on multiple nodes, each consisting of 8 NVIDIA H100 80 GB GPUs, interconnected via InfiniBand HDR. The software we use for training is the Megatron-LM library Shoeybi et al. (2019).
We use and modify the Megatron-LM (core v0.8.0) library for our experiments [^3] https://github.com/NVIDIA/Megatron-LM. Models are trained with data type bfloat16. Other optimization libraries used include FlashAttention Dao et al. (2022) and TransformerEngine [^4] https://github.com/NVIDIA/TransformerEngine. See the example scripts provided in supplementary material.
@@ B.2 Model Configuration
Let us elaborate more on our model configuration. The intermediate hidden dimension size, $d_{\rm MLP}$, is set to be four times the hidden dimension size, i.e., $4d_{\rm model}$. Bias is not used in the linear layers. We do not consider efficiency-motivated implementations like grouped query attention as well. Attention head number is chosen to increase with model size following standard practices. Other designs of the architecture follow Llama2’s closely (Touvron et al., 2023).
We vary model sizes keeping the ratio $n_{\rm layer}/d_{\rm model}$ to lie in the range 32 to 64, as in Kaplan et al. (2020). Smaller models for used for ablation studies. See Table 2 for the exact numbers for the model configuration.
@@ B.3 Training Configuration
As discussed in the main text, we adopt the WSD learning rate schedule for all experiments. The number of warmup steps is set approximately equal to the model size, as suggested by Porian et al. (2024). In the final phase of training, the learning rate decays linearly to 10% of its peak value, with the decay phase spanning roughly 10% of the total training steps, following the setup in Hägele et al. (2024). To emulate varying token budgets, we save intermediate checkpoints at logarithmically spaced intervals.
The general training configuration is summarized in Table 3, while model-specific hyperparameters, such as warmup iterations, initialization standard deviation ($\sqrt{2/5d_{\rm model}}$ (Le Scao et al., 2022)), maximum training steps, batch size, and tuned learning rate, are provided in Table 4. Batch size is scaled with model size according to standard practice, without tuning for optimality. The number of tokens is increased with model size as well, such that the parameter-to-token ratio remains roughly constant across different model sizes. We note that no specialized techniques for mitigating training instabilities, such as Z-loss or QK normalization, are employed, as such instabilities do not arise in our experiments.
@@
Table 4: Model-dependent training configuration. ”init. size” refers to the standard deviation of the normal distribution used for initializing the weights. ”Std iter.” refers to the number of iteration run on the model to reach about 70 times the model size in tokens, the standard training length used in the scaling law experiments. For several models, we also run longer training for exploration.
Model warmup iter. init. size Std iter. batch size LR
@@
B.4 On the choice of Data Domains Studied
Here, we argue that our experimental setup for CPT incorporates a significant domain shift, and it is highly relevant to frontier model research.
  • Domain gap in our setup. The base models are purposefully pretrained exclusively on the Common Crawl portion of the Slimpajama-DC dataset. This explicitly excludes data heavily skewed toward specialized domains such as mathematics (like arXiv) or code (like dedicated GitHub repositories), thereby maximizing the domain gap for the subsequent CPT stage.
  • Small overlap of domains. While some minimal token overlap between the generic web corpus and the specialized domains is inevitable, it is quantitatively small. We estimate that math/code data accounts for approximately 0.3 percent of the total Slimpajama-DC corpus.  [^5] Estimated based on filtering performed for the OpenCoder (Huang et al., 2025) dataset, which extracted math/code data from a similar CommonCrawl-derived corpus, FineWeb (Penedo et al., 2024). Consequently, the domain shift in our CPT experiments, which move to dedicated code and math corpora, is substantial.
  • Relevance to frontier LLM training. The strategy of pretraining mainly on generic web data and subsequently applying CPT on specialized domains like code (Stack/StarCoder) and mathematics (OpenWebMath) is a standard and highly relevant practice for state-of-the-art LLMs.
  • Model growth. For model growth experiments, the primary goal is to isolate and quantify the effect of increasing model capacity, not domain shift. Therefore, the second stage is intentionally conducted on the same generic web dataset as the base model, allowing us to accurately quantify the scaling behavior of the growth technique itself.
To summarize, our current findings are robust within the most practically relevant and challenging domain-transfer scenarios currently employed for large-scale LLM training.
B.5 GPU Hours and Costs
@@
The estimate is as follows. We calculate the FLOPs for training the largest first and second-stage models with maximum iterations using the $6ND$ approximation, ignoring the additional FLOPs required to continued pretrain models with shorter iterations (as we can reuse the intermediate checkpoints).  [^6] We have more than 450 runs; estimated from 25 data points for each of the 8 methods considered in Table 1, and 125 data points for the two joint scaling law experiments (CPT and model growth). We further assume that the per-second TFLOPs of the GPU is 400. We obtain around $10^{9}$ TFLOPs. Taking into account of additional experimentation and ablation runs, we estimate that around 3000 GPU hours were used for the entire project. Assuming a cost of 2 USD per GPU hour, the total cost is around 10,000 USD.
Scaling law methodology and cost justification. We note that our study contains over 450 total runs, as accurately mapping the relationship between multiple scaling variables (model size $N$, first-stage tokens $D_{1}$, and second-stage tokens $D_{2}$) necessitates a multitude of controlled experiments on models of various sizes. This methodology contrasts with single, large-scale training runs (e.g., training one 7B model from scratch), which would provide fewer data points for fitting complex scaling laws. This rigorous approach, balanced with the need for a reasonable overall budget, dictates our focus on models up to 1B parameters for extensive experimentation.
@@ C.1 Ablation of Learning Rate Schedules
Here, we compare the performances of using WSD and the commonly used learning rate (LR) cosine schedules (decaying to 10% of the peak LR value) (Touvron et al., 2023). The model size we use is 0.1B, with the following training configuration: batch size 512, 4,000 training iterations, and 200 warmup iterations. We can see from Figure 5 that both schedules yield similar (with WSD achieving slightly better final loss) performances. This justifies our choice of using the WSD schedule throughout the paper.
@@ C.3 Evaluation with Standard Benchmarks
We compare the performance of our trained 1B base model against existing models with similar sizes, Pythia (Biderman et al., 2023) and TinyLlama (Zhang et al., 2024b), based on standard natural language processing benchmarks, ARC (Clark et al., 2018), lambada (Paperno et al., 2016), logiqa (Liu et al., 2021), piqa (Bisk et al., 2020), sciq (Welbl et al., 2017), and winogrande (Sakaguchi et al., 2021).
@@
Table 5: Benchmarks’ performance comparison across models. Reported scores are accuracies (normalized by byte length whenever applicable). The first two columns are scores of existing models. The last column is evaluation results of the largest base model trained with the most number of tokens in this work. Our models are evaluated with the LM Evaluation Harness v0.4.0 library (Gao et al., 2024).
@@ C.4 Continual Pretraining Details
Using a lower LR for CPT is also a common practice (Ibrahim et al., 2024). As an ablation, we have experimented with using a constant LR schedule setting the value to be the final LR for the first-stage base model training, on a 0.1B model. We find that it yields worse performance (see Figure 7). Hence, we use the same LR (and LR schedule) as the base model for CPT.
@@ C.5 Model Growth Details
Stacking details. We first note that stacking does not preserve function but is empirically found to work well in practice (Du et al., 2024). We use the recommended stacking procedure in Du et al. (2024), which is, letting $M$ be the non-embedding part of the base model, the stacked model with growth factor $k$ is given by:
@@
Figure 7: model reuse with lower LR. We show that using the same LR as the base model achieves better final performance than using a lower LR for a 0.1B model continual pretrained on code data (left) and stacking (right).
@@ D.1 Fitted Exponents and Other Results
We show the fitted exponents of the multiplicative scaling laws studied in Table 1 in Table 6.
@@
Figure 8: $D_{1}$ has power-law scaling. We show scaling behavior of fist-stage training tokens ($D_{1}$) for different values of second-stage tokens ($D_{2}$), indicating that $D_{1}$ also has power-law scaling. Top left: Continual pretraining on code. Top right: Continual pretraining on mathematics data. Bottom left: Model growth by stacking (growth factor 2). Bottom right: Model growth by expansion (growth factor 2).
@@
D.2 Another variant of CPT scaling law.
When the dataset used for CPT is the same as the base model, we expect the following continuous scaling formula shown below holds:
$L(D_{1},D_{2})=A(D_{1}+D_{2})^{-\alpha}+E$ (11)
Hence, one may expect this holds for CPT on a different dataset as well, as assumed in Que et al. (2024); Wang et al. (2025). Fitting the above formula for the code (math) CPT scenario, we find that the RMS error is 0.0213 (0.0235), higher than the best ones in Table 1. Hence, our scaling law models CPT better, in contrast to previous assumptions, which may have overlooked the discontinuous (overtraining) effects of CPT.
One may further wonder if the model growth methods follow the above scaling law, since both stages use the same dataset. However, we find that the RMS error for (growth factor 2) expansion and stacking methods are 0.032 and 0.025 respectively, again larger than the ones in Table 1. Even though the same dataset is used in training, the model capacity is changed in model growth, explaining why the continuous scaling law does not hold. In Figure 9, we show that the continuous scaling law does not fit the data well visually.
D.3 Joint Scaling Law
@@
Figure 9: Left: Fits with different functional forms. We show in red the fits of the continuous scaling law does not capture the saturation effect well compared to our proposed multiplicative scaling law. Right: Fits of the joint upcycling scaling law of stacking. 10% of the collected data points with lowest losses are used for validation (orange points).
We provide the fitted coefficients of the joint scaling law in Table 7. In addition to model reuse, we also show the fitted coefficients for base models trained from scratch on the same dataset as the second stage (code data for CPT, and the same data as the base model for model growth). For model growth, we consider a growth factor of 2, and fit the parameters with model size $N$ before stacking for comparison conveniences. We further note that the fitted coefficients are produced by fitting the joint scaling law to all data points collected, including those used for validation in Figure 9.
@@
D.4 Why Fitting Scaling Laws Separately?
We justify our decision to fit two separate scaling laws—one for models trained from scratch and another for model reuse—instead of employing a unified formulation that spans all stages.
First, the scaling behavior of models trained from scratch is expected to differ from that of models undergoing second-stage pretraining. Specifically, at the limit $D_{1}\to 0$, grown models are initialized from a pretrained base model, whereas scratch-trained models start from random weights. This difference in initialization leads to distinct learning dynamics and, consequently, different scaling law parameters.
Similarly, the scaling behavior of the base model differs from that of the second-stage training in the limit $D_{2}\to 0$. As shown in Figure 7, second-stage training often begins with a rewarming phase, during which the loss initially increases before decreasing. This early instability deviates from the expected scaling of dense models, although the overall trend remains correlated—supporting the validity of Condition 1, due to function preservation and empirical observations that rewarming does not entirely disrupt loss behavior (as quoted from Clauset et al. (2009): “In practice, few empirical phenomena obey power laws for all values of xx”).
@@ D.4 Why Fitting Scaling Laws Separately?
Appendix E More Observations and Implications of the Scaling Laws
E.1 Implication based on Scaling Laws’ Extrapolation
We consider an implication based on the extrapolation of the scaling laws, motivated by studies on MoE upcycling (Komatsuzaki et al., ; Liew et al., 2025).
A key practical consideration is how much of the initial investment in pretraining, the so-called sunk cost ($D_{1}$), can be effectively leveraged in the second stage of training. This question has been explored in the context of MoE upcycling strategies. For example, it was found that upcycling yields benefits up to 120% of the sunk cost. That is, to match the performance of an upcycled MoE model that underwent an additional 0.4 trillion tokens of training after an initial 2T tokens, training a comparable MoE from scratch would require 2.4T tokens—representing an effective saving of 2T tokens.
We investigate this in model growth by stacking (growth factor 2). We define $D^{*}$ as the number of tokens required for training from scratch to match the performance of a grown model with the same sunk cost, following Liew et al. (2025):
$L^{\rm scratch}_{2N}(D^{*})=L^{\rm grown}_{N}(D_{1}=D^{*},D_{2}=D^{*})$ (12)
Figure 10: Model growth efficiency decreases with sunk cost and model size. For token budgets above the curve(s), training from scratch is more efficient than stacking-based model growth (growth factor 2); for budgets below, model growth remains advantageous. Shown are the numerical (blue) and analytical (red) solutions of Equation 12. Vertical axis is $D^{*}$ as in Equation 12.
E.2 Compute Optimality
For scaling studies of LLMs, the cost of training is often estimated using floating point operations (FLOPs), which we simply refer to as compute. The compute cost can be approximated as $C=6ND$ (Kaplan et al., 2020). One is often interested in training LLMs with the least amount of compute. For CPT, optimizing the compute based on Equation 6 leads to the scaling relations:
where $\alpha_{\rm eff}\coloneq\alpha_{2}-\alpha_{3}\log D_{1}$. Notably, as $D_{1}$ increases, $\alpha_{\rm eff}$ decreases, meaning larger models require more tokens for compute-optimal model reuse.
@@ E.2 Compute Optimality
Notably, as $D_{1}$ increases, $\alpha_{\rm eff}$ decreases, meaning larger models require more tokens for compute-optimal model reuse. We leave empirical verification of these relationships to future work, primarily due to cost considerations (see Appendix B.5 for an estimate of GPU hours spent in the work).
E.3 Training with Extremely Large second-stage Token-to-Parameter Ratio
15M-to-30M model. We show the loss values of the last three $D_{2}$ points (each run with different $D_{1}$’s) in Table 8. To estimate the stochastic noise on the loss, for each $D_{2}$, we train again from the nearest stable-phase checkpoint but with three different random seeds. Although there is slight indication of crossover (comparing $D_{1}=0.2884,9.2275$ at $D_{2}=9.2275,27.6824$), we are unable to make strong claims on it without a better understanding of the noise effects.
0.5B-to-1B model. We also train the larger 0.5B-to-1B model to more extreme token-to-parameter ratios (up to 200). The result is shown in Figure 11 (and Table 9 for the last four measured $D_{2}$), where we do not see significant crossover at the large $D_{2}$ limit.
Furthermore, at large $D_{2}$, the points are slightly curved upwards. This can best be seen by modifying the scaling law using a similar argument given in Section 4.2, by adding a new fitting parameter $D_{\rm 2,max}$:
$\hat{D_{2}}=\left(D_{2}^{-1}+D_{\rm 2,max}^{-1}\right)^{-1},$ (14)
The mean square error of the ”tail” for the original and new scaling law are 2.7e-5 and 1.3e-5 (with similar global mean square errors) indicating that the new one fits the tail better (and from the Figure the new scaling law has an upward curve at large $D_{2}$). This new formulation predicts a crossover and saturation of the curves afterwards (crossover occurs only when the fitted $D_{\rm 2,max}\lesssim e^{\alpha_{2}/\alpha_{3}}$). The predicted crossover is beyond the computational reach and is in the region where noises are dominant.
Figure 11: Stacking with extremely large token-to-parameter ratio (0.5B-to-1B). We compare the original scaling law and the new scaling law with a new fitting parameter $D_{\rm 2,max}$.
Table 8: 15M-to-30M models’ cross-entropy losses and the estimated error from multiple runs with different random seeds.
$D_{2}$ (B) $D_{1}$ (B) Loss Error
9.2275 0.2884 3.069959 $\pm$ 0.001345
0.4325 3.069911
0.5767 3.078497
1.1534 3.068825
2.3069 3.070050
9.2275 3.073884
18.4549 0.2884 3.055554 $\pm$ 0.004026
1.1534 3.065518
2.3069 3.060454
9.2275 3.063905
27.6824 0.2884 3.055849 $\pm$ 0.005699
1.1534 3.053557
2.3069 3.051366
9.2275 3.054023
Table 9: 0.5B-to-1B models’ cross-entropy losses. The estimated error is 0.0012-0.0016.
$D_{2}$ (B) $D_{1}$ (B) Loss
147.6395 9.2275 2.199658
18.4549 2.198132
36.9099 2.195471
73.8198 2.191318
184.5494 9.2275 2.182314
36.9099 2.179196
73.8198 2.176350
221.4593 9.2275 2.178030
36.9099 2.175260
73.8198 2.172848
258.3691 9.2275 2.165875
36.9099 2.163731
73.8198 2.161603
258.3691 9.2275 2.165875
36.9099 2.163731
73.8198 2.161603
E.4 More on Mechanistic Explanations
We show the gradient norms of model growth by stacking in Figure 12, where we also see that the gradient norms decrease with $D_{1}$ for fixed $D_{2}$, indicating saturation effects.
Figure 12: Overtrained models have larger gradient norms when undergoing model reuse. We show the gradient norm curves for model growth (0.5B to 1B model via stacking) with respect to number of training steps in the second stage for base models trained with different values of first-stage tokens ($D_{1}$).
Next, we study other possible mechanistic explanations of the saturation effects. Our scope of investigation is inspired by Lyle et al. (2023, 2024); Klein et al. (2024), which studied mechanistic signatures of plasticity loss in continual learning.
Parameter norms.
Increased parameter norms (or weight norms) are indicated to be a signature of plasticity loss (Nikishin et al., 2022). However, as in Figure 13, we do not observe a clear trend of parameter norms with $D_{1}$ (which correlates directly with saturation) when training our models. See the caption of Figure 13 for more details.
Figure 13: Parameter norms do not show clear trend with $D_{1}$. We show the parameter norm curves training a 0.1B base model with respect to number of training steps, where the end of each line corresponds to the parameter norm of a base model trained with the respective number of training steps, $D_{1}$. The parameter norms of models (of various $D_{1}$) we use increase before decreasing with respect to $D_{1}$ (i.e., non-monotonic). There is therefore no direct correlation with $D_{1}$ (which correlates directly with saturation).
Sharpness via training trajectory.
Lyle et al. (2023) suggested to look at the variance of training curves as a proxy for sharpness, i.e., increased variance in the training curves indicates that it is harder for the model to navigate the sharper loss landscape, indicating plasticity loss. As in Figure 14, we do not observe a clear trend of difference in variance of training curves for different values of $D_{1}$ either (except in early training where training with smaller $D_{1}$ has slightly lower variance). See the caption of Figure 14 for more details.
Figure 14: Variance of training curves do not show clear trend with $D_{1}$. We show the variance of training curves for model growth by stacking (0.5B to 1B) with respect to number of training steps in the second stage for base models trained with different values of first-stage tokens ($D_{1}$). Up to 10,000 steps, the variance is slightly smaller for the training curve with lower $D_{1}$. However, there is no clear trend of variance in the training curves for different $D_{1}$ onwards.
Sharpness via Hessian.
Another way of studying the loss landscape is through the top eigenvalue of the Hessian ($\lambda_{max}$). Here, we present a preliminary analysis using this quantity. We use the Hessian-vector-product and power iteration to estimate $\lambda_{max}$ on training samples using base models trained with different $D_{1}$. In Figure 15, we show the estimated $\lambda_{max}$ with respect to $D_{1}$.
Figure 15: Hessian with respect to $D_{1}$. We show the Hessian ($\lambda_{max}$) of a 0.1B model base models trained with different values of first-stage tokens ($D_{1}$), on model reuse scanarios (CPT on code, math; stacking) and single-stage training scenario. The number of power iteration in calculating $\lambda_{max}$ is 10, and the number of samples used to test each scenario is 128.
The overall trend is not very clear, but roughly speaking, we see that the model reuse scenarios (CPT on code, math; stacking) have some correlation with $D_{1}$, while training with the same data does not seem to correlate with $D_{1}$.
To summarize, we only observe a clear trend of gradient norms with $D_{1}$ among the mechanistic quantities studied. We leave the studies of other possible mechanisms (and a more detailed study of the Hessian) to future work.
E.5 More on the Toy Model
Theoretical derivation of the exponential distribution w.r.t. depth. The exponential decay of probability with respect to depth $d$ is a derivation from the maximum entropy principle motivated by Mandelbrot’s work on the distribution of word frequencies (Mandelbrot and others, 1953).
We seek a probability distribution $P(d)$ that maximizes Shannon Entropy $H=-\sum P(d)\ln P(d)$ subject to normalization $\sum P(d)=1$ and a fixed average cost $\sum P(d)C(d)=\bar{C}$. Using Lagrange multipliers:
$L=-\sum P(d)\ln P(d)-\lambda_{0}\left(\sum P(d)-1\right)-\lambda\left(\sum P(d)C(d)-\bar{C}\right)$
Setting $\frac{\partial L}{\partial P(d)}=0$ solving for $P(d)$ yields the exponential (Boltzmann) distribution:
$P(d)\propto e^{-\lambda\cdot C(d)}$
In our tree model, depth $d$ represents the search cost $C$. This decay results in the power-law distribution of features.
More on Hutter’s learning curve theory. Let us describe the setup of Hutter (2021) in more detail: Each datum is associated with a feature $k$, and the feature is memorized when it is observed during training. Assuming that i.i.d. data are drawn from a distribution $P(k)$, the expected loss is then proportional to the probability of encountering an unobserved feature during training after $D$ training samples:
$L\propto\sum_{k=1}^{\infty}P(k)\left(1-P(k)\right)^{D}$
If $P(k)$ follows a power-law distribution $P(k)\sim k^{-(1+\alpha)}$, the loss can be shown to scale as $L\sim D^{-\alpha/(1+\alpha)}\approx D^{-\alpha}$ for small $\alpha$.
Note that a simpler alternative assumption that leads to the same conclusion at small $\alpha$ is that features at the tail of the feature distribution is simply not observed during training. The loss is then an integral over the tail:
$L=\int_{D}^{\infty}k^{-(1+\alpha)}\textrm{d}k\sim D^{-\alpha},$
leading to the same scaling behavior.